Two cylindrical containers are shown. Container A has radius r and height h and holds a maximum of 24 ounces of water. Container B has radius 3r and height 3h.
What is the maximum amount of water that Container B can hold?

Answer by stanbon(75887)

Two cylindrical containers are shown. Container A has radius r and height h and holds a maximum of 24 ounces of water. Container B has radius 3r and height 3h.
What is the maximum amount of water that Container B can hold?
-----
Volume of smaller container = pi*r^2*h
----
Volume of larger container = pi(3r)^2(3h) = 27
------
Water in smaller = 24 oz
Water in larger = 27*24 = 648 oz
-----------
Cheers,
Stan H.